Normal conversations at 50cm correspond to a sound level recorded at 65db. What is the new sound level if two other people join in?
Answer #1

We know that intensity level = 10*log(I/i0) where I is intensity of a persons voice

when 2 more people join, intensity becomes 3 times,

intensity level = 10*log(3*I/i0) = 10*log 3 + 10*log(I/i0) = 4.8 + 65 dB

= 69.8 dB answer